OpenStudy (anonymous):

Use a system of equations to solve the following problem: Fairview High School has a total of 92 boys and girls who play sports. If the number of boys is 17 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

girls = x boys = 2x+17 3x+17= 92 >> 3x=75 >> x=25 boys= 2x+17=2(25)+17 = 67 boys